# Study on quantum thermalization from thermal initial states in a superconducting quantum computer

### v2

@@ Abstract
Quantum thermalization in contemporary quantum devices, in particular quantum computers, has recently attracted significant theoretical interest. Unusual thermalization processes, such as the Quantum Mpemba Effect (QME), have been explored theoretically. However, there is a shortage of experimental results due to the difficulty in preparing thermal states. In this paper, we propose a method to indirectly address this challenge. Moreover, we experimentally validate our approach using IBM quantum devices, providing results for unusual relaxation in equidistant quenches as predicted for the IBM qubit. We also assess the formalism introduced for the QME, obtaining results consistent with the theoretical predictions. This demonstration underscores that our method can streamline the investigation of thermal states and thermalization in quantum physics.
@@ I Introduction
The current Noisy Intermediate-Scale Quantum (NISQ) era quantum computers are highly susceptible to decoherence [1]. This poses a current limitation on the potential for promised exponential speedup of quantum algorithms over their classical counterparts. However, as will be demonstrated in this work, NISQ era devices serve as a controllable platform for studying and testing Open Quantum Systems (OQS) physics [2]. In particular, we will illustrate that these devices can be employed for the study of quantum thermalization. Thermal states hold significant theoretical interest; for instance, the QME [3], an analog to the classical phenomenon where hot water freezes faster than cold water [4], has been predicted for certain quantum systems [5, 6]. However, experimentally preparing thermal states is challenging as it would necessitate precise control of the environmental temperature. While some algorithms allow the initialization of a thermal state [7], utilizing them for testing arbitrary states is impractical.
@@
In this work, we introduce a method for indirectly studying the physics of initial thermal states in any quantum system, provided that we can prepare the eigenstates of the Hamiltonian. We validate this method using two IBM quantum computers, experimentally testing the formalism for the QME in this system. Furthermore, we examine another unusal thermalization process known as Equidistant Quenches [8, 9], which contrasts the relaxation into equilibrium of a pure state with that of an initially equidistant thermal state.
@@
The IBM Quantum Experience [10] provides public access to superconducting qubits [11]. These devices undergo a relaxation process into a thermal equilibrium state due to interaction with their environment, leading to a loss of information. To solve this problem, quantum error correction [12] can be applied. However, such techniques require several additional qubits for correcting errors on a given qubit, making them impractical given that the current size of the largest quantum computers is on the order of $\sim$ 100 qubits. Alternatively, pulse sequences can be applied to protect the state from decoherence; for example, Dynamical Decoupling has been tested on IBM devices [13]. Another approach involves designing the quantum algorithm to be inherently immune to external noise [14].
@@
The paper is structured as follows: In Section 2, we introduce the method and present the theoretical results essential for this work. Section 3 outlines the methodology employed in the demonstrations. The results obtained from the IBM devices are presented in Section 4, followed by the conclusion in Section 5.
II Theoretical Background
@@
This equation has an analytical solution derived in [8]. In the same article, the concept of Equidistant Quenches was introduced, where, under certain conditions, a thermal state and a pure state initially at the same distance from equilibrium relax differently. For a system following Equation 5, there is a critical value for the dephasing rate, $\gamma_{z,c}$:
@@
In the case of a system described by Equation 5 the effect should not appear since we always have that $\lambda_{3}=0$ and $\lambda_{2}<0$ [8].
III IBMQ Demonstrations
@@
IV Results
For thermal states on ibmq_manila, in Figure 2 and Figure 3 it is shown that, for both distance functions, we observe a monotonic increase with temperature at all times. This indicates that we do not observe the QME, which is consistent with the theory that does not predict QME to occur in a system described by Equation 5.
@@ IV Results
In Figures 4 and 5, we compare the relaxation of two different initial thermal states with both distance functions. Once again, there is no crossing between the curves. The equilibrium temperature found was $T_{eq}=9.5$ mK.
Figure 4: Trace distance as a function of time for two different initial thermal states on ibmq_manila, $T_{cold}=20$ mK and $T_{hot}=30$ mK.
Figure 5: (a) Entropic distance as a function of time for two different initial thermal states on ibmq_manila, $T_{cold}=20$ mK and $T_{hot}=30$ mK.
@@
As observed, the equidistant thermal state does depend on the distance function, illustrating the fact that temperature in quantum systems is ill-defined. The equilibrium state was determined by fitting the time evolution of |$\ket{0}$, providing an estimate for the computer temperature of $T_{eq}=10.3$ mK.
V Discussion
In this work, we have introduced and successfully tested a powerful method for exploring the physics of thermal states in quantum systems, which, in principle, could be extended to other quantum systems. This significantly simplifies the challenge of obtaining experimental results in the field of quantum thermodynamics. More significantly, we can demonstrate quantum thermalization from thermal initial states directly. We have also provided experimental evidence of coherent states relaxing faster into equilibrium than equidistant thermal states, as predicted in [8] and [9] theoretically. Furthermore, we have identified strong non-Markovian effects on the IBMQ devices, in line with [16], offering additional experimental insights into this matter. These findings are crucial to consider when implementing decoherence control techniques [13]. We observed that thermal states on the quantum computer follow a normal, Markovian relaxation into equilibrium, suggesting that Markovian equations may be sufficient for describing this type of states in this particular system under longer time limit.
However, it is important to highlight the limitations and weaknesses of the proposed method. Firstly, despite the formal equivalence between eigenstates and thermal state evolution, thermal states are not actually prepared on the quantum system. Therefore, our method is not applicable for quantum algorithms that utilize thermal states as a computational tool. Additionally, ensuring identical interactions with the environment for each eigenstate evolution is necessary. We have addressed this by running the different circuits as closely in time as possible. Further investigation of this method on different quantum systems with higher complexity is needed.
@@ Acknowledgements.
We would like to thank Dr Dawei Luo for discussions. We gratefully acknowledge supports from the Basque Country Government (Grant No. IT1470- 22) and Grant No. PGC2018-101355-B-I00 funded by MCIN/AEI/10.13039/501100011033.

### v3

@@ Abstract
Quantum thermalization in contemporary quantum devices, in particular quantum computers, has recently attracted significant theoretical interest. Unusual thermalization processes, such as the Quantum Mpemba Effect (QME), have been explored theoretically. However, there is a shortage of experimental results due to the difficulty in preparing thermal states. In this paper, we propose a protocol to indirectly address this challenge. Moreover, we experimentally validate our protocol using IBM quantum devices, providing results for unusual relaxation in equidistant quenches as predicted for the IBM qubit. We also assess the formalism introduced for the QME, obtaining results consistent with the theoretical predictions. This demonstration underscores that our protocol can provide an alternative way of studying thermal states physics when their direct preparation may be too difficult.
@@ I Introduction
The current Noisy Intermediate-Scale Quantum (NISQ) era quantum computers are highly susceptible to decoherence [1]. This poses a current limitation on the potential for promised exponential speedup of quantum algorithms over their classical counterparts. However, as will be demonstrated in this work, NISQ era devices serve as a controllable platform for studying and testing Open Quantum Systems (OQS) physics [2]. In particular, we will illustrate that these devices can be employed for the study of quantum thermalization. Thermal states hold significant theoretical interest; for instance, the QME [3], an analog to the classical phenomenon where hot water freezes faster than cold water [4], has been predicted for certain quantum systems [5, 6].
However, since thermal states are mixed states, they cannot be obtained from a pure initial state via unitary evolution. Preparing initial thermal states directly from experimental thermalization is highly impractical and would require precise control of the environment temperature at all times. Another possibile approach is to utilize quantum algorithms that prepare thermal states [7, 8, 9, 10, 11]. This algorithms typically rely on variational quantum algorithms (VQA’s) and have been tested on NISQ devices [8, 12]. This approach is of interest and utility for many quantum computation applications that require thermal state inializations, such as quantum simulation [13] or quantum machine learning [14]. Nonetheless, the study of quantum thermodynamics necessitates examining a large quantity of thermal states and the current algorithms have applicability and scalability issues [9]. Thus, the experimental cost of preparing many individual thermal states and manipulating them for the intended application is enormous. In this work, we address this issue by introducing a protocol for indirectly studying the afterward dynamics of initialized thermal states in any quantum system, provided that we can prepare pure eigenstates of a Hamiltonian. While our proposal cannot resolve the challenge of thermal state preparation, we will demonstrate that it significantly simplifies the task of studying the dynamics of initial thermal states. Therefore, it holds great practiality for basic research applications.
@@
We apply this protocol using two IBM quantum computers, experimentally testing the formalism for the QME in this system. Furthermore, we examine another unusal thermalization process known as Equidistant Quenches [15, 16], which contrasts the relaxation into equilibrium of a pure state with that of an initially equidistant thermal state.
@@
The IBM Quantum Experience [17] provides public access to superconducting qubits [18]. These devices undergo a relaxation process into a thermal equilibrium state due to interaction with their environment, leading to a loss of information. To solve this problem, quantum error correction [19] can be applied. However, such techniques require several additional qubits for correcting errors on a given qubit, making them impractical given that the current size of the largest quantum computers is on the order of $\sim$ 100 qubits. Alternatively, pulse sequences can be applied to protect the state from decoherence; for example, Dynamical Decoupling has been tested on IBM devices [20]. Another approach involves designing the quantum algorithm to be inherently immune to external noise [21, 22].
@@
The paper is structured as follows: In Section 2, we introduce the protocol and present the theoretical results essential for this work. Section 3 outlines the methodology employed in the demonstrations. The results obtained from the IBM devices are presented in Section 4, followed by the conclusion in Section 5.
@@
Introducing this Hamiltonian in equation 5, the time evolution has an analytical solution that can be writen as:
$\rho_{00}=$ $(\rho_{00}(0)-\rho_{00}^{eq})e^{-t\gamma(e^{\beta\hbar\omega}+1)}+\rho_{00}^{eq},$ (7)
$\rho_{01}=$ $\rho_{01}(0)e^{-\frac{1}{2}t(\gamma+e^{-\beta\hbar\omega}\gamma)}e^{-4\gamma_{z}t}e^{-i\omega t},$
$\rho_{10}=$ $\rho_{10}(0)e^{-\frac{1}{2}t(\gamma+e^{-\beta\hbar\omega}\gamma)}e^{-4\gamma_{z}t}e^{i\omega t},$
$\rho_{11}=$ $(\rho_{11}(0)-\rho_{11}^{eq})e^{-t\gamma(e^{\beta\hbar\omega}+1)}+\rho_{11}^{eq}.$
Where $\rho_{ij}$ are the reduced density matrix components and $\rho_{ij}^{eq}$ are the equilibrium state components. The derivation can be found in [15]. In this same article, the concept of Equidistant Quenches was introduced, where, under certain conditions, a thermal state and a pure state initially at the same distance from equilibrium relax differently. For a system following Equation 5, there is a critical value for the dephasing rate, $\gamma_{z,c}$:
@@
In the case of a system described by Equation (5) the effect should not appear since we always have that $\lambda_{3}=0$ and $\lambda_{2}<0$ [15]. Thus, for the IBM superconducting qubit, normal thermalization as defined in [3] is expected for the initial thermal states. That is, the distance to equilibrium should be a monotonically increasing function of temperature at all times for $T>T_{eq}$.
II Methods
@@
Table 1: Calibration values at the time of the demonstration.
The mixture probabilities used for constructing the thermal states are calculated using Equations (3) and (6) along with the values from Table 1. We present them in Figure 2.
Figure 2: Thermal state eigenstates probabilities used for the thermal state construction as a function of temperature for (a) ibmq_belem and (b) ibmq_manila.
Communication with IBM services was performed using Qiskit. Data treatment was done in Python, utilizing the Matplotlib library for plotting and the Scipy package for numerical fitting of Equation 7. We left the equilibrium component as a free parameter and used the calibration values given in Table I for the calculation of the dephasing and relaxation rates.
III Results
For thermal states on ibmq_manila, in Figure 3 and Figure 4 it is shown that, for both distance functions, we observe a monotonic increase with temperature at all times. This indicates that we do not observe the QME, which is consistent with the theory that does not predict QME to occur in a system described by Equation (5).
@@ III Results
In Figures 5 and 6, we compare the relaxation of two different initial thermal states with both distance functions. Once again, there is no crossing between the curves. The equilibrium temperature found was $T_{eq}=58$ mK.
Figure 5: Trace distance as a function of time for two different initial thermal states on ibmq_manila, $T_{cold}=80$ mK and $T_{hot}=90$ mK.
Figure 6: (a) Entropic distance as a function of time for two different initial thermal states on ibmq_manila, $T_{cold}=80$ mK and $T_{hot}=90$ mK.
@@
As observed, the equidistant thermal state does depend on the distance function, illustrating the fact that temperature in quantum systems is ill-defined. The equilibrium state was determined by fitting the time evolution of |$\ket{0}$, providing an estimate for the computer temperature of $T_{eq}=67$ mK.
IV Discussion
In this work, we have introduced and successfully tested a powerful protocol for exploring the physics of thermal states in quantum systems, which, in principle, could be extended to other quantum systems. This significantly simplifies the challenge of obtaining experimental results in the field of quantum thermodynamics. More significantly, we can demonstrate quantum thermalization from thermal initial states directly. We have also provided experimental evidence of coherent states relaxing faster into equilibrium than equidistant thermal states, as predicted in [15] and [16] theoretically. Furthermore, we have identified strong non-Markovian effects on the IBMQ devices, in line with [24], offering additional experimental insights into this matter. These findings are crucial to consider when implementing decoherence control techniques [20]. We observed that thermal states on the quantum computer follow a normal, Markovian relaxation into equilibrium, suggesting that Markovian equations may be sufficient for describing this type of states in this particular system under longer time limit.
However, it is important to highlight the limitations and weaknesses of the proposed protocol. Firstly, despite the formal equivalence between eigenstates and thermal state evolution, thermal states are not actually prepared on the quantum system. Therefore, our protocol may not be applicable for quantum algorithms that utilize thermal states as a computational tool. Additionally, ensuring identical interactions with the environment for each eigenstate evolution may be necessary. We have addressed this by running the different circuits as closely in time as possible. Moreover, this protocol needs to be further studied on different quantum systems with higher complexity. In particular, it would be desirable to apply our protocol on a highly controlled environment in order to further validate the protocol.
@@ Acknowledgements.
We would like to thank Dr Dawei Luo for discussions. We gratefully acknowledge supports from the Basque Country Government (Grant No. IT1470- 22) and Grant No. PGC2018-101355-B-I00 funded by MCIN/AEI/10.13039/501100011033. This project has also received support from the Spanish Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call - Quantum Spain, EU through the Recovery, Transformation and Resilience Plan – NextGenerationEU within the framework of the Digital Spain 2026.